Circular No. 6562

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1997 D1 Jean Mueller reports her discovery of a comet on a plate taken by K. M.

Rykoski and J. D. Mendenhall with the 1.2-m Oschin Schmidt telescope in the course of the second Palomar Sky Survey. The comet is slightly diffuse with some central condensation; no tail is visible. The following positions from beginnings and ends of trails were reduced by B. G. Marsden, following measurement by J. Mueller:

1997 UT  α_{2000}  δ_{2000}  m_{1}
Feb. 17.46667  12^{h}48^{m}21.^{s}29  +46^{o}23′44′′.1  16
17.50833  12 48 15.70  +46 24 34.6
19.51736  12 43 12.93  +47 08 50.0
19.53472  12 43 10.81  +47 09 07.3

COMET P/1997 C1 (GEHRELS) New orbital elements for this comet (IAUC 6549, 6550), published on

MPEC 1997-D06 and MPC 29067 (and used for the ephemeris below), show it to be of short period (P = 17 years).

1997 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Feb. 21 9^{h}01.^{m}02 +13^{o}07.^{′}8 3.430 4.377 161.^{o}1 4.^{o}2 17.1
Mar. 3 8 56.72 +13 30.8 3.522 4.409 150.1  6.4 17.2
13 8 53.59 +13 49.9 3.640 4.442 139.3  8.4 17.3
23 8 51.86 +14 04.0 3.780 4.475 128.9 10.0 17.4
Apr. 2 8 51.59 +14 12.5 3.938 4.508 119.0 11.2 17.5
12 8 52.77 +14 15.2 4.108 4.541 109.4 12.0 17.6
22 8 55.32 +14 12.0 4.287 4.574 100.3 12.5 17.8

COMET 118P/SHOEMAKER-LEVY 4 Total visual magnitude estimates: Jan. 14.98 UT, 12.8 (R. J. Bouma,

Groningen, The Netherlands, 0.25-m reflector); Feb. 2.75, 12.8 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); 8.89, 12.5 (J. Carvajal, Madrid, Spain, 0.44-m reflector); 14.79, 12.6 (M. Plsek, Lelekovice, Czech Republic, 0.35-m reflector).

SUPERNOVA 1996cb IN NGC 3510

On IAUC 6524, line 3, for SN 1996ca read SN 1996cb

1997 February 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green